# Key Words.: Galaxies: clusters: general – Galaxies: clusters: individual: Abell 426 – X-rays: galaxies: clusters – Radio continuum: galaxies – Galaxies: clusters: intracluster medium – Galaxies: kinematics and dynamics1 Introduction

### v1

@@
Perseus cluster in its X-ray entirety with SRG/eROSITA et al. I. Merger and Radio-Uroboroses.
The Perseus cluster (Abell 426) is a nearby massive galaxy cluster that spans several degrees. We combined SRG/eROSITA, XMM-Newton, and Chandra data to get a complete coverage of this cluster in X-rays up to and beyond, although at the largest radii, spatial non-uniformities of X-ray background and foreground dominate. While the Perseus central part represents a canonical cool-core structure with clear signs of AGN Feedback, the outskirts, in turn, serve as a convincing example of a merger-perturbed system. X-ray data suggest that IC310 is the main galaxy of a subcluster that merges with Perseus over the past . Overall, this configuration resembles the merger between the Coma cluster and the NGC4839 group. It is statistically more likely to find a merging group near the apocenter of its orbit. Therefore, it is not surprising that IC310 in Perseus has a relatively small velocity relative to the main cluster, similarly to NGC4839 in Coma.
@@
The Perseus cluster (Abell 426) is the X-ray brightest and one of the best-studied clusters. Its core properties were used to develop the concept of ”AGN feedback” (e.g., Boehringer et al., 1993; Churazov et al., 2000; Fabian et al., 2003; Zhuravleva et al., 2015). In the core and at larger scales, X-ray images show signs of perturbations, including a sequence of sloshing patterns located at different distances from the center (e.g., Branduardi-Raymont et al., 1981; Schwarz et al., 1992; Ettori et al., 1998; Churazov et al., 2003; Fabian et al., 2011; Simionescu et al., 2012; Tamura et al., 2014; Zhang et al., 2020; Sanders et al., 2020; Walker et al., 2022; Zhu et al., 2021), (see, also Markevitch & Vikhlinin, 2007; Zuhone & Roediger, 2016, for reviews on merger-induced perturbations).
@@ 2 X-ray data
For X-ray analysis, we use the data accumulated by the eROSITA (Predehl et al., 2021) telescope onboard the SRG X-ray observatory (Sunyaev et al., 2021) in the course of its all-sky survey. Initial reduction and processing of the data were performed using standard routines of the eSASS software (Brunner et al., 2018; Predehl et al., 2021), while the imaging and spectral analysis were carried out with the background models, vignetting, point spread function (PSF) and spectral response function calibrations built upon the standard ones via slight modifications motivated by results of calibration and performance verification observations (e.g. Churazov et al., 2021; Khabibullin et al., 2023).
@@ 2.1 Large-scale image
A 10x10 degrees patch of the sky centered at the Perseus cluster is shown in Fig. 1. The eROSITA 0.4-2.3 keV image illustrates the connection of Perseus to the large-scale structure, in particular, two bright clusters 2A0251+413 () and CIZAJ0300.7+4427 (), which trace filamentary structures in the vicinity of the main cluster. The same structures are visible in the PLANCK -parameter map (Fig. 1, middle panel), where is proportional to the line-of-sight integrated electron pressure and can be derived from distortions of the Cosmic Microwave Background (see Zeldovich & Sunyaev, 1969, where the formulae describing these distortions were found).
The location of the Perseus cluster close to the Galactic Plane introduces some biases in the X-ray and -maps. The extinction derived from the Bayestar data (Green et al., 2019) varies from 0.1 to 3.6 across the image. As a result, significant variations of photoelectric absorption modulate the X-ray image, and the dust-correlated structures are also visible in the -map (see, e.g., the bottom-left corner of the images). Assuming that can be converted to equivalent hydrogen column density as (e.g., Predehl & Schmitt, 1995; Güver & Özel, 2009) and considering the contribution of atomic gas (traced by HI, ), we can expect a modulation of the Perseus X-ray flux in the 0.4-2.3 keV up to within a circle of 1.4 degrees. This estimate is done for the APEC spectrum with a temperature of 5 keV. The total attenuation factor of the cluster flux (relative to ) is about 0.55.
@@
It is useful to compare the X-ray images with the distribution of galaxies with redshifts comparable to that of the cluster. To this end, we used the 2MASS Redshift Survey (Huchra et al., 2012) and selected galaxies with line-of-sight recession velocity in the range from to . The distribution of these galaxies is shown in Fig. 2. The left (20x20 degrees) image shows clearly the famous large-scale filaments in the vicinity of Perseus, which are part of the Perseus-Pisces supercluster. The most prominent two filaments, relevant for this study, are marked with black lines. One of them is nearly horizontal in equatorial coordinates and is dominated by galaxies with redshifts similar to the main cluster. Another one is diagonal in this plot, and its upper end is dominated by galaxies with larger recession velocities (blue colors). The right image (2x2 degrees) is the zoomed version of the same plot. The chain of galaxies between NGC1275 and IC310 appears to align with the former filament, while NGC1265 might be part of the latter.
@@ 2.3 X-ray composite image
Focusing on the X-ray emission inside the Perseus virial radius, we would like to complement the eROSITA coverage of the entire field with higher angular resolution of Chandra and XMM-Newton in the Perseus central region. We first generated separate background-subtracted, exposure and vignetting-corrected X-ray images for eROSITA (0.4-2.3 keV; 4” pixels), XMM-Newton (0.5-3.5 keV; 2” pixels), and Chandra (0.5-3.5 keV, 1” pixels). These images have been combined in two steps. In the first step, the mean surface brightness in a circle centered at NGC1275 has been calculated, and the normalization of images was adjusted to compensate for slight differences in the energy bands and effective areas. The resulting images were rebinned to 2” resolution and co-added using a position-dependent weighting factor. This factor favors the Chandra images in the core and the eROSITA image at the largest distances from NGC1275. The resulting 2”-resolution, degrees image is shown in Fig. 3. The inset shows the central 1% of the same image, where Chandra’s resolution captures familiar structures of the X-ray emission in the Perseus core.
@@ 2.4 Departures from a symmetric model
The known elongation of the Perseus X-ray emission along the East-West direction is already visible in Fig. 3. To emphasize further the departures from spherical symmetry, we fitted the simplest beta model to the radial X-ray surface brightness profile. The brightest compact sources in the field were excised prior to fitting the model. The profile and the best-fitting beta model are shown in the Appendix B. The dynamic range covered by the profile from the inner to approaches a factor of . Overall, the beta model gives a reasonable (for our purposes) approximation of the Perseus X-ray surface brightness. We note in passing that at radii in the range 1-3 degrees, a stray light can give a non-negligible or even dominant contribution to the surface brightness profile. However, to reveal asymmetric features in the outskirts of the Perseus cluster, this is not required as long as the flux from the cluster core dominates the stray light and, therefore, the model remains symmetric, even though the amplitude of deviations can be affected.
Fig. 4 shows the X-ray image divided by the model. The model is the sum of the cluster emission (the -model described above) and the constant sky background. Once again, the core shows the familiar pattern of X-ray cavities and several spiral-like structures, while on larger scales, the dominant features are two large regions of excess emission on both sides of the core, which are roughly aligned with the East-West direction. Many of these structures (or parts of them) have already been seen in previous X-ray observations (e.g., Branduardi-Raymont et al., 1981; Schwarz et al., 1992; Ettori et al., 1998; Churazov et al., 2003; Fabian et al., 2011; Simionescu et al., 2012; Tamura et al., 2014; Walker et al., 2022; Zhu et al., 2021), but the complete and uniform coverage of eROSITA reveals this pattern in its entirety.
@@ 3.1 Merger geometry and timing
The dynamic state of the Perseus cluster (outside the very core) is a subject of many studies over several decades, (see, e.g., Branduardi-Raymont et al., 1981; Kent & Sargent, 1983; Andreon, 1994; Churazov et al., 2003; Fabian et al., 2011; Simionescu et al., 2012; Bellomi et al., 2024; Sanders et al., 2020; Kang et al., 2024; HyeongHan et al., 2025, among others) with an overall conclusion that the cluster is neither fully relaxed nor in a state of a major merger.
@@ 3.3.1 IC310
As mentioned above, IC310 appears to be at the center of the subcluster/group that merges with the main cluster. In terms of IC310 offset position ( from the core) and low line-of-sight velocity ( relative to NGC1275)555Quoted relative line-of-sight velocities are calculated as ., this galaxy is reminiscent of the NGC4839 group in the Coma cluster. The low velocity might be related to the proximity of the group to the apocenter, hence low 3D and line-of-sight velocities. While IC310 is originally classified as a narrow-tail galaxy, which would be more consistent with a high relative velocity of the galaxy relative to ICM (Begelman et al., 1979), it might be a special case of a radio galaxy seen nearly edge-on Sijbring & de Bruyn (1998); Gendron-Marsolais et al. (2020). This interpretation is supported by the blazar-like properties of the nucleus (e.g. Kadler et al., 2012; Aleksić et al., 2014), with a pc-scale jet making a relatively small angle to the line of sight666Another interesting detail discussed by Kadler et al. (2012) is that the orientation of the pc-scale jet is similar to that of the kpc-long tail.. The most recent LOFAR data revealed a much more extended radio tail of IC310 (van Weeren et al., 2024). The tail makes a sharp turn towards the Perseus center (see Fig. 7). This morphology could be explained by a merger scenario, where IC310 has already gone through the pericenter and the apocenter and is now moving back towards the core. In this case, the ”reversal” in the far tail direction is explained by the so-called slingshot effect (Lyskova et al., 2019; Sheardown et al., 2019; Zhang et al., 2019), when the ICM shifted backwards by the ram pressure, swings to the opposite side when the motion of the halo slows down, and then reverses near the apocenter. This configuration leads to a loop-like geometry of the tail that we dubbed ”radio-Uroboros”.
@@ 3.3.2 NGC1265
Unlike IC310, NGC1265 has a very high velocity relative to NGC1275 () and it lacks obvious large-scale X-ray emission around the X-ray bright core. While there are even higher velocity galaxies in the same region (plausibly associated with another large-scale filament, see Fig. 2), the radio tail suggests interaction with the Perseus ICM. The tail identified in Ryle & Windram (1968); Miley et al. (1975), led to the classification of NGC1265 as a Narrow-angle-tail galaxy, although at higher resolution, the core rather looks like wide-angle-tail galaxy (see a discussion of tags and boxes in Rudnick, 2021). The extent of its long and bent tail (projected length ) was revealed with WSRT observations (Sijbring & de Bruyn, 1998) and more recently imaged in great detail (Gendron-Marsolais et al., 2020; van Weeren et al., 2024).
Remarkably, despite being very different from IC310, the radio tail of NGC1265 also qualifies for the ”radio-Uroboros” tag (see Fig. 7). The statistic of radio tails’ direction suggests that for a bare galaxy, especially moving with a very high velocity, the gas feeding the central AGN can be stripped during the first infall (e.g., Roberts & Parker, 2020; Lal et al., 2022), NGC1265 appears to be a newcomer to the Perseus cluster. Yet, it has a very extended tail (probably a factor of a few longer than its apparent projected size). In Pfrommer & Jones (2011), the curious properties of the NGC1265 tail are explained as due to the galaxy crossing an accretion shock near the virial radius, with the galaxy located at the near side of the cluster. While this is a plausible scenario, we note that the ICM in the Perseus can be significantly perturbed by the ongoing merger, and for a galaxy on a nearly straight trajectory, dictated by its high velocity, the ICM motions can perturb the radio tail left by NGC1265. We, therefore, consider below a simple model with NGC1265 at the far side (along the line of sight) of the main cluster. This model broadly corresponds to scenarios c) and d) in Sijbring & de Bruyn (1998).
@@ 3.3.3 Test particle (NGC1265) in a static potential
As the starting point, we consider a test particle moving in a static potential well having an NFW mass distribution (Navarro et al., 1997). We further assume a zero total energy for this particle, i.e, . We chose a coordinate system with and in the sky plane, and along the line-of-sight. We aim to reproduce the line-of-sight velocity and the projected distance of NGC1265. The requirement to have zero total energy implies that the 3D distance of the particle from the cluster center is less than , while the lower limit is set by the observed projected distance from the cluster center, i.e., (see Fig. 8). If the marginally bound particle is close to , it has to move exactly along the line of sight. In the opposite case, the angle to the line of sight is set by the relation between the local 3D velocity (set by the 3D radius) and the . For the limiting case of , this angle is about degrees. For any realistic 3D position (further away than ), this angle will be smaller. If the overall shift of the tail eastwards of the current position of NGC1265 reflects the large-scale trajectory of the galaxy, it should be crossing the main cluster on a slightly curved trajectory. Given the high velocity of NGC1265, the curved trajectory requires the galaxy to cross the core of the cluster. A sample of such trajectories is shown in Fig. 9. They all correspond to the same (current) projected distance, line-of-sight velocity, and the velocity direction in the sky plane and differ in the magnitude of the 3D velocity. This is, of course, a small subsample of possible trajectories (given direct observational constraints and our assumption that the galaxy is currently on the far side of the cluster). In this model, we assume that gas motions in the region between NGC1275 and IC310 further disturb the tail trajectory to give it the observed shape. No attempt was made to get constraints on the detailed properties of the tail, given that the density and pressure of the ICM change significantly along the galaxy trajectory, and the power of the NGC1265 jets can vary too.
The above model does not consider constraints coming from the uncertainties associated with the Faraday Rotation measurement in the direction of NGC1265 (see, e.g., de Bruyn & Brentjens, 2005; Brentjens, 2011), that might predominantly come from the Milky Way rather than from the cluster (see Pfrommer & Jones, 2011, for a scenario with NGC1265 on the near side of the cluster).

### v2

# Perseus cluster in its X-ray entirety with SRG/eROSITA

@@
Perseus cluster in its X-ray entirety with SRG/eROSITA
The Perseus cluster (Abell 426) is a nearby massive galaxy cluster that spans several degrees. We combined SRG/eROSITA, XMM-Newton, and Chandra data to get a complete coverage of this cluster in X-rays up to and beyond, although at the largest radii, spatial non-uniformities of the X-ray sky background and foreground dominate. While the Perseus central part represents a canonical cool-core structure with clear signs of AGN Feedback, the outskirts, in turn, serve as a convincing example of a merger-perturbed system. X-ray data suggest that IC310 is the main galaxy of a subcluster that merges with Perseus over the past . Overall, this configuration resembles the merger between the Coma cluster and the NGC4839 group. It is statistically more likely to find a merging group near the apocenter of its orbit. Therefore, it is not surprising that IC310 in Perseus has a relatively small velocity relative to the main cluster, similarly to NGC4839 in Coma.
@@
The Perseus cluster (Abell 426) is the X-ray brightest and one of the best-studied clusters. Its core properties were used to develop the concept of ”AGN feedback” (e.g., Boehringer et al., 1993; Churazov et al., 2000; Fabian et al., 2003; Zhuravleva et al., 2015). In the core and at larger scales, X-ray images show signs of perturbations, including a sequence of sloshing patterns located at different distances from the center (e.g., Branduardi-Raymont et al., 1981; Schwarz et al., 1992; Ettori et al., 1998; Churazov et al., 2003; Fabian et al., 2011; Simionescu et al., 2012; Tamura et al., 2014; Zhang et al., 2020; Sanders et al., 2020; Walker et al., 2022; Zhu et al., 2021), (see, also Markevitch and Vikhlinin, 2007; Zuhone and Roediger, 2016, for reviews on merger-induced perturbations).
@@ 2 X-ray data
For X-ray analysis, we use the data accumulated by the eROSITA (Predehl et al., 2021) telescope onboard the SRG X-ray observatory (Sunyaev et al., 2021) in the course of its all-sky survey. Initial reduction and processing of the data were performed using standard routines of the eSASS software (Brunner et al., 2018; Predehl et al., 2021), while the imaging analysis was carried out with the detector background models, vignetting, point spread function (PSF) and spectral response function calibrations built upon the standard ones via slight modifications motivated by results of calibration and performance verification observations (e.g. Churazov et al., 2021a; Khabibullin et al., 2023). These modifications are intended to reproduce best the effective (i.e., averaged over multiple source transits of the telescope’s Field-of-View during scanning mode observations) characteristics, including the averaged energy-dependent effective area and extended wings of the PSF (see AppendixA for details). Given the very bright central core of the Perseus cluster, the effect of the stray light scattering, included in the extended PSF model, might be particularly relevant for the radial profile measurement beyond of the cluster. At these radii, the emission of the cluster becomes comparable or smaller than the sum of the astrophysical and instrumental background. The eROSITA’s detector background is well measured and was very stable during most of the sky survey after exclusion of periods of enhanced Solar activity (see, e.g., Predehl et al., 2021; Sunyaev et al., 2021; Freyberg et al., 2021).
@@ 2.1 Large-scale image
A 10x10 degrees patch of the sky centered at the Perseus cluster is shown in Fig. 1. The eROSITA 0.4-2.3 keV image illustrates the location of the Perseus cluster with respect to the large-scale structure traced by massive halos. In particular, two bright clusters AWM7/2A0251+413 () and CIZAJ0300.7+4427 () are clearly visible. AWM7 together with Perseus belongs to the Perseus-Pisces supercluster (e.g., Böhringer et al., 2021) and is well aligned with a major filament going through Perseus in the East-West direction. CIZAJ0300.7+4427 has a larger recession velocity, and the direct connection to Perseus is less certain. The same clusters are also visible in the PLANCK -parameter map (Fig. 1, middle panel), where is proportional to the line-of-sight integrated electron pressure and can be derived from distortions of the Cosmic Microwave Background (see Zeldovich and Sunyaev, 1969, where the formulae describing these distortions were found).
The location of the Perseus cluster close to the Galactic Plane introduces some biases in the X-ray and -maps. The extinction derived from the Bayestar data (Green et al., 2019) varies from 0.1 to 3.6 across the image. As a result, significant variations of photoelectric absorption modulate the X-ray image, and the dust-correlated structures are also visible in the -map (see, e.g., the bottom-left corner of the images). Assuming that can be converted to equivalent hydrogen column density as (e.g., Predehl and Schmitt, 1995; Güver and Özel, 2009) and considering the contribution of atomic gas (traced by HI, ), we can expect a modulation of the Perseus X-ray flux in the 0.4-2.3 keV up to within a circle of 1.4 degrees. This estimate is done for the APEC spectrum with a temperature of 5 keV. The total attenuation factor of the cluster flux (relative to ) is about 0.55. We further assess the impact of this modulation in Appendix D.
@@
It is useful to compare the X-ray images with the distribution of galaxies with redshifts comparable to that of the cluster. To this end, we used the 2MASS Redshift Survey (Huchra et al., 2012) and selected galaxies with line-of-sight recession velocity in the range from to . The distribution of these galaxies is shown in Fig. 2. The left (20x20 degrees) image shows clearly the famous large-scale filaments in the vicinity of Perseus, which are part of the Perseus-Pisces supercluster. The two most prominent filaments, relevant for this study, are marked with black lines. One of them is nearly horizontal in equatorial coordinates and is dominated by galaxies with redshifts similar to the main cluster. Another one is diagonal in this plot, and its upper end is dominated by galaxies with larger recession velocities (blue colors). The right image (2x2 degrees) is the zoomed version of the same plot. The chain of galaxies between NGC1275 and IC310 appears to align with the former filament, while NGC1265 might be part of the latter.
@@ 2.3 X-ray composite image
Focusing on the X-ray emission inside the Perseus virial radius, we would like to complement the eROSITA coverage of the entire field with the higher angular resolution of Chandra and XMM-Newton in the Perseus central region. We first generated separate detector-background-subtracted, exposure and vignetting-corrected X-ray images for eROSITA (0.4-2.3 keV; 4” pixels), XMM-Newton (0.5-3.5 keV; 2” pixels), and Chandra (0.5-3.5 keV, 1” pixels). These images have been combined in two steps. In the first step, the mean surface brightness in a circle centered at NGC1275 has been calculated, and the normalization of images was adjusted to compensate for slight differences in the energy bands and effective areas. The resulting images were rebinned to 2” resolution and co-added using a position-dependent weighting factor. This factor favors the Chandra images in the core and the eROSITA image at the largest distances from NGC1275. The resulting 2”-resolution, degrees image is shown in Fig. 3. The inset shows the central 1% of the same image, where Chandra’s resolution captures familiar structures of the X-ray emission in the Perseus core.
@@ 2.4 Departures from a symmetric model
The known elongation of the Perseus X-ray emission along the East-West direction is already visible in Fig. 3. To emphasize further the departures from spherical symmetry, we fitted the simplest beta model to the radial X-ray surface brightness profile. Here, is the spatially constant sky X-ray background. The brightest compact sources in the field were excised prior to fitting the model. The profile and the best-fitting beta model are shown in the Appendix C. The dynamic range covered by the profile from the inner to approaches a factor of . Overall, the beta model gives a reasonable (for our purposes) approximation of the Perseus X-ray surface brightness. We note in passing that at radii in the range 1-3 degrees, a stray light can give a non-negligible or even dominant contribution to the surface brightness profile. However, to reveal asymmetric features in the outskirts of the Perseus cluster, this is not required as long as the flux from the cluster core dominates the stray light and, therefore, the model remains symmetric, even though the amplitude of deviations can be affected.
Fig. 4 shows the X-ray image divided by the model, specifically, . Once again, the core shows the familiar pattern of X-ray cavities and several spiral-like structures, while on larger scales, the dominant features are two large regions of excess emission on both sides of the core, which are roughly aligned with the East-West direction. Many of these structures (or parts of them) have already been seen in previous X-ray observations (e.g., Branduardi-Raymont et al., 1981; Schwarz et al., 1992; Ettori et al., 1998; Churazov et al., 2003; Fabian et al., 2011; Simionescu et al., 2012; Tamura et al., 2014; Walker et al., 2022; Zhu et al., 2021), but the complete and uniform coverage of eROSITA reveals this pattern in its entirety.
@@ 3.1 Merger geometry and timing
The dynamic state of the Perseus cluster (outside the very core) is a subject of many studies over several decades, (see, e.g., Branduardi-Raymont et al., 1981; Kent and Sargent, 1983; Andreon, 1994; Churazov et al., 2003; Fabian et al., 2011; Simionescu et al., 2012; Bellomi et al., 2024; Sanders et al., 2020; Kang et al., 2024; HyeongHan et al., 2025, among others) with an overall conclusion that the cluster is neither fully relaxed nor in a state of a major merger.
@@ 3.3.1 IC310
As mentioned above, IC310 appears to be at the center of the subcluster/group that merges with the main cluster. In terms of IC310 offset position ( from the core) and low line-of-sight velocity ( relative to NGC1275)555Quoted relative line-of-sight velocities are calculated as ., this galaxy is reminiscent of the NGC4839 group in the Coma cluster. The low velocity might be related to the proximity of the group to the apocenter, hence low 3D and line-of-sight velocities. While IC310 is originally classified as a narrow-tail galaxy, which would be more consistent with a high relative velocity of the galaxy relative to ICM (Begelman et al., 1979), it might be a special case of a radio galaxy seen nearly edge-on Sijbring and de Bruyn (1998); Gendron-Marsolais et al. (2020). This interpretation is supported by the blazar-like properties of the nucleus (e.g. Kadler et al., 2012; Aleksić et al., 2014), with a pc-scale jet making a relatively small angle to the line of sight666Another interesting detail discussed by Kadler et al. (2012) is that the orientation of the pc-scale jet is similar to that of the kpc-long tail.. The most recent LOFAR data revealed a much more extended radio tail of IC310 (van Weeren et al., 2024). The tail makes a sharp turn towards the Perseus center (see Fig. 7). This morphology could be explained by a merger scenario, where IC310 has already gone through the pericenter and the apocenter and is now moving back towards the core. In this case, the ”reversal” in the far tail direction is explained by the so-called slingshot effect (Lyskova et al., 2019; Sheardown et al., 2019; Zhang et al., 2019), when the ICM shifted backwards by the ram pressure, swings to the opposite side when the motion of the halo slows down, and then reverses near the apocenter. This configuration leads to a loop-like geometry of the tail that we dubbed ”radio-Uroboros”.
@@ 3.3.2 NGC1265
Unlike IC310, NGC1265 has a very high velocity relative to NGC1275 () and it lacks obvious large-scale X-ray emission around the X-ray bright core. While there are even higher velocity galaxies in the same region (plausibly associated with another large-scale filament, see Fig. 2), the radio tail suggests interaction with the Perseus ICM. The tail identified in Ryle and Windram (1968); Miley et al. (1975), led to the classification of NGC1265 as a Narrow-angle-tail galaxy, although at higher resolution, the core rather looks like a wide-angle-tail galaxy (see a discussion of tags and boxes in Rudnick, 2021). The extent of its long and bent tail (projected length ) was revealed with WSRT observations (Sijbring and de Bruyn, 1998) and more recently imaged in great detail (Gendron-Marsolais et al., 2020; van Weeren et al., 2024).
Remarkably, despite being very different from IC310, the radio tail of NGC1265 also qualifies for the ”radio-Uroboros” tag (see Fig. 7). The statistics of radio tails’ direction suggest that for a bare galaxy, especially moving with a very high velocity, the gas feeding the central AGN can be stripped during the first infall (e.g., Roberts and Parker, 2020; Lal et al., 2022), NGC1265 appears to be a newcomer to the Perseus cluster. Yet, it has a very extended tail (probably a factor of a few longer than its apparent projected size). In Pfrommer and Jones (2011), the curious properties of the NGC1265 tail are explained as due to the galaxy crossing an accretion shock near the virial radius, with the galaxy located at the near side of the cluster. While this is a plausible scenario, we note that the ICM in the Perseus can be significantly perturbed by the ongoing merger, and for a galaxy on a nearly straight trajectory, dictated by its high velocity, the ICM motions can perturb the radio tail left by NGC1265. We, therefore, consider below a simple model with NGC1265 at the far side (along the line of sight) of the main cluster. This model broadly corresponds to scenarios c) and d) in Sijbring and de Bruyn (1998).
@@ 3.3.3 Test particle (NGC1265) in a static potential
As the starting point, we consider a test particle moving in a static potential well having an NFW mass distribution (Navarro et al., 1997). We further assume a zero total energy for this particle, i.e, . We chose a coordinate system with and in the sky plane, and along the line-of-sight. We aim to reproduce the line-of-sight velocity and the projected distance of NGC1265. The requirement to have zero total energy implies that the 3D distance of the particle from the cluster center is less than , while the lower limit is set by the observed projected distance from the cluster center, i.e., (see Fig. 8). If the marginally bound particle is close to , it has to move exactly along the line of sight. In the opposite case, the angle to the line of sight is set by the relation between the local 3D velocity (set by the 3D radius) and the . For the limiting case of , this angle is about degrees. For any realistic 3D position (further away than ), this angle will be smaller. If the overall shift of the tail eastwards of the current position of NGC1265 reflects the large-scale trajectory of the galaxy, it should be crossing the main cluster on a slightly curved trajectory. Given the high velocity of NGC1265, the curved trajectory requires the galaxy to cross the core of the cluster. A sample of such trajectories is shown in Fig. 9. They all correspond to the same (current) projected distance, line-of-sight velocity, and the velocity direction in the sky plane, and differ in the magnitude of the 3D velocity. This is, of course, a small subsample of possible trajectories (given direct observational constraints and our assumption that the galaxy is currently on the far side of the cluster). In this model, we assume that gas motions in the region between NGC1275 and IC310 further disturb the tail trajectory to give it the observed shape. No attempt was made to get constraints on the detailed properties of the tail, given that the density and pressure of the ICM change significantly along the galaxy trajectory, and the power of the NGC1265 jets can vary too.
The above model does not consider constraints coming from the uncertainties associated with the Faraday Rotation measurement in the direction of NGC1265 (see, e.g., de Bruyn and Brentjens, 2005; Brentjens, 2011), that might predominantly come from the Milky Way rather than from the cluster (see Pfrommer and Jones, 2011, for a scenario with NGC1265 on the near side of the cluster).